# ECHO: Frequency-aware Hierarchical Encoding for Variable-length Signals

### v3

@@ 5.3 Experimental Results
1) Effect of dataset scaling. By comparing BEATs (71.86%), CED (71.28%) and EAT (74.23%) with Dasheng (76.04%) with base scale, we observe that incorporating additional large-scale datasets might help mitigate the domain mismatch between general audio pre-training and machine sound analysis. A similar trend can also be found in FISHER (76.86%) and our proposed ECHO (77.65%), both trained on additional audio datasets, showing that scale-up of training data consistently enhances cross-domain representation learning.
2) Sliding-patch vs. conventional patch modeling. Traditional foundation models (BEATs, CED, and EAT) rely on conventional patch tokenization, yielding total average scores in the range of 70–74% in our SIREN benchmark. In contrast, Dasheng introduces a sliding-patch strategy, leading to a higher overall mean of 76.04%, outperforming EAT (74.23%) by $1.81\%$. A similar observation is confirmed within the band-splitting family: our proposed ECHO reaches a total average of 77.65% on SIREN, surpassing FISHER (76.86%) by $0.79\%$. These results highlight that sliding-patch modeling is more effective than fixed patch partitioning for machine sound analysis.

### v4

@@ 5.3 Experimental Results
1) Effect of dataset scaling. By comparing BEATs (71.86%), CED (71.28%) and EAT (74.15%) with Dasheng (76.04%) with base scale, we observe that incorporating additional large-scale datasets might help mitigate the domain mismatch between general audio pre-training and machine sound analysis. A similar trend can also be found in FISHER (76.86%) and our proposed ECHO (77.65%), both trained on additional audio datasets, showing that scale-up of training data consistently enhances cross-domain representation learning.
2) Sliding-patch vs. conventional patch modeling. Traditional foundation models (BEATs, CED, and EAT) rely on conventional patch tokenization, yielding total average scores in the range of 70–74% in our SIREN benchmark. In contrast, Dasheng introduces a sliding-patch strategy, leading to a higher overall mean of 76.04%, outperforming EAT (74.15%) by $1.89\%$. A similar observation is confirmed within the band-splitting family: our proposed ECHO reaches a total average of 77.65% on SIREN, surpassing FISHER (76.86%) by $0.79\%$. These results highlight that sliding-patch modeling is more effective than fixed patch partitioning for machine sound analysis.